OpenStudy (anonymous):

Can someone verify my answer? multiply the following and combine terms where possible. -q(p+q) = -q - qp

jimthompson5910 (jim_thompson5910):

-q(p+q) = -qp - q*q = -qp-q^2 So you were close, but your first term should have been -q^2

OpenStudy (anonymous):

yeah I forgot to add that LOL